# Power $\Sigma_{1}$ in Card with two Woodin cardinals

### v1

@@ Abstract
This is a simplification of the version above. ∎
@@
Now $N|\delta_{0}^{N}=M|\delta_{0}^{N}$. For let $\theta<\delta_{0}^{N}$ be a cardinal and suppose $N|\theta=M|\theta$, and let $P\triangleleft M$ with $\rho_{\omega}^{P}=\theta$. Then by hypothesis (and since $\delta_{0}^{N}\leq\delta_{0}^{M}$), $T_{P}$ is wellfounded, so by Claims 6 and 7, $P\triangleleft N$, as desired.
Now since $N|\delta_{0}^{N}=M|\delta_{0}^{M}$ and $M\subseteq N$, it follows that $\delta_{0}^{M}$ is Woodin in $M$. So $\delta_{0}^{N}=\delta_{0}^{M}$, and so $\delta_{0}^{+N}=\delta_{0}^{+M}$.
A similar calculation as before now gives that $N|\delta_{1}^{N}=M|\delta_{1}^{N}$, and then again since $M\subseteq N$, it follows that $\delta_{1}^{N}=\delta_{1}^{M}$ and $\delta_{1}^{+N}=\delta_{1}^{+M}$, as desired. ∎

### v2

@@ 4. $L_{\varepsilon}[N]$ is a premouse modelling $\delta$ is Woodin”, and
If $P$ is iterable above $\theta$, then any $\bar{P}$ (as built by $T_{P}$) is also iterable above $\bar{\theta}=\varrho^{-1}(\theta)$, which implies that $T_{P}$ is wellfounded, much as before. Conversely, suppose that $\theta<\delta_{0}^{N}$, but $P$ is not above-$\theta$, $(\omega,\theta^{+}+1)$-iterable above $\theta$. Then we can fix an iteration tree witnessing this, together with the canonical Q-structures (or $M_{1}^{\#}(M(\mathcal{T}))$ satisfying $\delta(\mathcal{T})$ is Woodin”), since $N|\delta_{0}^{N}$ is sufficiently closed to compute these (as $\theta<\delta_{0}^{N}$). But then we can take a countable hull of these objects to produce a branch through $T_{P}$. ∎
Now fix $\theta,P$ again as above, and suppose that $P\models$ “there is a Woodin cardinal $\leq\theta$”. Then we define a tree $T_{P}^{\prime}$, similar to $T_{P}$, but now the Q-structures involved are just of form $\mathcal{J}_{\alpha}(M(\mathcal{T}))$ for some ordinal $\alpha$, and the special case of producing $M_{1}^{\#}(M(\mathcal{T}))$ satisfying $\delta(\mathcal{T})$ is Woodin” is replaced by $M(\mathcal{T})^{\#}\models$$\delta(\mathcal{T})$ is Woodin”.
Claim 8.
Let $\theta,P,T^{\prime}_{P}$ be as in the previous paragraph. Then:
  1. 1.
    If $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable then $T^{\prime}_{P}$ is wellfounded.
  2. 2.
    If $\theta<\delta_{1}^{N}$ and $T^{\prime}_{P}$ is wellfounded then $P$ is above-$\theta$, $(\omega,\theta^{+}+1)$-iterable.
Proof.
This is a simplification of the proof of Claim 7. ∎
@@
Now $M|\delta_{0}^{M}=N|\delta_{0}^{N}$. For let $\theta<\delta_{0}^{N}$ be a cardinal and suppose $M|\theta=N|\theta$, and let $P\triangleleft M$ with $\rho_{\omega}^{P}=\theta$. Then by hypothesis (and since $\delta_{0}^{N}\leq\delta_{0}^{M}$), $T_{P}$ is wellfounded, so by Claims 6 and 7, $P\triangleleft N$, as desired. Since also $\theta^{+M}=\theta^{+N}$, it follows that $M|\theta^{+M}=N|\theta^{+N}$.
Now since $M|\delta_{0}^{N}=N|\delta_{0}^{N}$ and $M\subseteq N$, it follows that $\delta_{0}^{M}$ is Woodin in $M$. So $\delta_{0}^{M}=\delta_{0}^{N}$, and so $\delta_{0}^{+M}=\delta_{0}^{+N}$.
A similar calculation as before now gives that $M|\delta_{1}^{N}=N|\delta_{1}^{N}$, and then again since $M\subseteq N$, it follows that $\delta_{1}^{M}=\delta_{1}^{N}$ and $\delta_{1}^{+M}=\delta_{1}^{+N}$, as desired. ∎